OpenStudy (anonymous):

Write the equation of the line given in the graph. -6,3,6

OpenStudy (anonymous):

not enough info in your question to provide an answer.

OpenStudy (anonymous):

they are the points

OpenStudy (annon):

it that x,y,z?

OpenStudy (anonymous):

theres a straight line which point is on -6 a point that lands on the 3 and than 6

OpenStudy (anonymous):

shenita a point in the x-y coordinate system has an x and a y like (1,1). A point in the xyz coordinate system will have 3 numbers like (1,1,1). Are you in the x-y coordinate plane or the xyz coordinate plane?

OpenStudy (annon):

x or y components?

OpenStudy (anonymous):

-6 is x 3 is y and 6 is x

OpenStudy (anonymous):

A point on the x / y axes consist of a pair of numbers, one for x and one for y. It takes two points to define a line. So you should be listing the (x,y) coordinates of two points.

OpenStudy (anonymous):

Ok so the points then are (-6,0), (0,3) and (6,0).

OpenStudy (anonymous):

(-6,3) (3,6)

OpenStudy (annon):

y=3x+3

OpenStudy (anonymous):

yes i guess thats how i ca explain the graph peer

OpenStudy (annon):

get it?

OpenStudy (anonymous):

no annon

OpenStudy (anonymous):

If it's (-6,3) and (3,6), then you determine the slope by calculating change in y over change in x: (6-3)/(3--6) = 3/9 = 1/3. That means for every 3 units in x you move, the line rises one unit in y. So calculating from the (-6,3) point, we move 6 in x and up 2 in y to find the y-intercept at (0,5). Using the formula y = mx + b, where m is slope and b is y-intercept, we have the answer: y = 1/3x + 5

OpenStudy (anonymous):

thanx for that ..im not good with graphs at all and dont know how to use my graphin calc..geesh its bee a while

OpenStudy (anonymous):

shenita is the graph a straight line or an upside-down parabola?

OpenStudy (anonymous):

straight <-------------------------> like that

OpenStudy (anonymous):

with points

OpenStudy (anonymous):

Ahh ok. Thanks.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

hickninja is right. Step 1 is to find the slope using this equation: m=(y2-y1)/(x2-x1). Then you have to use this equation: y=mx+b and plug in the slope (m) and one of your points to solve for the b. I used the point (-6,3) to plug into y=mx+b so that will give you b=5. Then the equation is y=(1/3)x +5

OpenStudy (anonymous):

ok